Dustin is 11 years younger than Elias. In two years, Dustin will be half as old as Elias. How old is Dustin?
Answer by checkley77(12844)

D=E-11
D+2=(E+2)/2
E-11+2=(E+2)/2
E-9=(E+2)2
2(E-9)=E+2
2E-18=E+2
2E-E=2+18
E=20 FOR ELIAS'S AGE.
D=20-11=9 FOR DUSTIN'S AGE.
PROOF:
9+2=(20+2)/2
11=22/2
11=11